# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v3

@@ B.7 Results on UDA Benchmarks
CDTrans [66] DEiT IN 97.1 90.5 82.4 77.5 96.6 96.1 93.6 88.6 97.9 86.9 90.3 62.8 88.4
PMTrans [75] ViT-B IN-21k 98.9 93.7 84.5 73.3 99.0 98.0 96.2 67.8 94.2 98.4 96.6 49.0 87.5
TVT [67] ViT-B IN-21k 97.1 92.9 85.3 66.4 97.1 97.1 89.3 75.5 95.0 94.7 94.5 55.1 86.7
TVT* [67] ViT-B IN-21k 95.8 85.8 81.9 68.4 95.9 96.2 91.9 70.3 93.8 93.7 92.9 48.5 84.6
TVT [67] ViT-B DinoV2 98.9 88.7 90.3 64.2 99.3 74.5 95.3 66.0 85.3 94.6 97.9 54.6 84.1
TVT [67] ViT-B Ours 97.0 89.9 89.4 73.8 98.0 88.9 94.4 85.9 93.8 94.5 97.7 54.3 88.2
SSRT [60] ViT-B IN-21k 98.9 87.6 89.1 84.8 98.3 98.7 96.3 81.1 94.9 97.9 94.5 43.1 88.8
SSRT [60] ViT-B DinoV2 99.2 88.1 89.9 85.4 98.4 98.9 97.6 84.8 96.2 97.1 95.4 48.3 89.9
SSRT [60] ViT-B Ours 99.4 88.6 91.4 87.9 98.3 99.1 97.1 88.0 95.9 98.1 96.0 51.2 90.9
@@ B.8 Attention Map Visualizations
To aid our analysis in section 7, we visualize attention scores for different ViT blocks across multiple models, including DinoV2, D-[L]-$r64$ (i.e. the last row of table 3), the second and third rows of table 3, MAE, SatMAE, and M-[L]-$r64$. These visualizations are shown in fig. 9 for 3 different images from the validation set of fMoW-RGB. Since our models are trained without registers, we truncate attention scores more than 5 standard deviations away from the mean, thus removing artifact attention scores with unusually high values on background patches (Darcet et al., 2023).
@@ Appendix C Training Details
In this section, we describe hyperparameters and hardware configurations used for our models.
@@ C.2 PEFT Fine-Tuning
For data augmentations, we only use the drop-path augmentation (Larsson et al., 2016) at a rate of 0.2, with no dropout, mixup, or cutmix. We note that the original LoRA configuration outperforms other PEFT techniques when paired with the drop-path regularization technique. For example, we find that BOFT does not pair well with drop-path, instead performing most effectively with a custom multiplicative dropout technique (Liu et al., 2023). We include the result with the best hyperparameter configuration for each row in table 1.
@@ C.4 Multi-Spectral Images
C.5 Downstream datasets
@@
C.6 Dataset Licenses
@@ Appendix D Environmental Impact
Following (Cong et al., 2022), we compare the carbon footprint of pre-training using ExPLoRA with domain-specific solutions such as SatMAE. We use the carbon footprint calculator proposed by Lacoste et al. (2019). Our results are in table 18.

### v4

@@ B.7 Results on UDA Benchmarks
1    CDTrans [70] DEiT IN 97.1 90.5 82.4 77.5 96.6 96.1 93.6 88.6 97.9 86.9 90.3 62.8 88.4
2    PMTrans [80] ViT-B IN-21k 98.9 93.7 84.5 73.3 99.0 98.0 96.2 67.8 94.2 98.4 96.6 49.0 87.5
3    TVT [71] ViT-B IN-21k 97.1 92.9 85.3 66.4 97.1 97.1 89.3 75.5 95.0 94.7 94.5 55.1 86.7
4    TVT* [71] ViT-B IN-21k 95.8 85.8 81.9 68.4 95.9 96.2 91.9 70.3 93.8 93.7 92.9 48.5 84.6
5    TVT [71] ViT-B DinoV2 98.9 88.7 90.3 64.2 99.3 74.5 95.3 66.0 85.3 94.6 97.9 54.6 84.1
6    TVT [71] ViT-B Ours 97.0 89.9 89.4 73.8 98.0 88.9 94.4 85.9 93.8 94.5 97.7 54.3 88.2
7    SSRT [62] ViT-B IN-21k 98.9 87.6 89.1 84.8 98.3 98.7 96.3 81.1 94.9 97.9 94.5 43.1 88.8
8    SSRT [62] ViT-B DinoV2 99.2 88.1 89.9 85.4 98.4 98.9 97.6 84.8 96.2 97.1 95.4 48.3 89.9
9    SSRT [62] ViT-B Ours 99.4 88.6 91.4 87.9 98.3 99.1 97.1 88.0 95.9 98.1 96.0 51.2 90.9
@@ B.8 Attention Map Visualizations
To aid our analysis in section 7, we visualize attention scores for different ViT blocks across multiple models, including DinoV2, D-[L]-$r64$ (i.e. the last row of table 3), the second and third rows of table 3, MAE, SatMAE, and M-[L]-$r64$. These visualizations are shown in fig. 9 for 3 different images from the validation set of fMoW-RGB. Since our models are trained without registers, we truncate attention scores more than 5 standard deviations away from the mean, thus removing artifact attention scores with unusually high values on background patches (Darcet et al., 2024).
@@ Appendix C Training Details
In this section, we describe hyperparameters and hardware configurations used for our models. We also include details on datasets used for our experiments.
@@ C.2 PEFT Fine-Tuning
For data augmentations, we only use the drop-path augmentation (Larsson et al., 2016) at a rate of 0.2, with no dropout, mixup, or cutmix. We note that the original LoRA configuration outperforms other PEFT techniques when paired with the drop-path regularization technique. For example, we find that BOFT does not pair well with drop-path, instead performing most effectively with a custom multiplicative dropout technique (Liu et al., 2024). We include the result with the best hyperparameter configuration for each row in table 1.
@@ C.4 Multi-Spectral Images
C.5 Dataset Information
We include detailed information about the datasets used in our experiments. Table 18 shows the number of samples in the train and validation splits for all datasets used in this work.
@@ C.5 Dataset Information
Since each dataset contains a varying number of training images, the number of ExPLoRA pre-training iterations should be adjusted accordingly. We include the recommended number of ExPLoRA pre-training iterations in table 18. Note, however, that this number can be varied depending on other hyperparameters such as batch size, learning rate, LoRA rank, and number of unfrozen blocks.
In general, we find that ExPLoRA demonstrates the largest performance gains on larger datasets where sufficient diversity prevents overfitting during extended pre-training. For datasets with fewer than 50k training samples (e.g., SpaceNet V1, GlobalWheat), the improvements from ExPLoRA are more modest, as the limited data diversity can lead to overfitting with extensive pre-training. Conversely, ExPLoRA works exceptionally well on larger datasets such as fMoW-(RGB, Sentinel, Temporal), Camelyon17, and VisDA2017, where 100k-200k pre-training iterations provide substantial performance improvements while maintaining computational efficiency compared to full domain-specific pre-training.
@@
Dataset #Train #Validation ExPLoRA Iters Domain
fMoW-RGB 363.6k 53.0k 200k Remote Sensing
fMoW-Sentinel 712.9k 84.9k 80k Remote Sensing (Multi-spectral)
fMoW-Temporal 83.4k 14.2k 80k Remote Sensing (Temporal)
SpaceNet V1 6.0k 1.5k 10k Remote Sensing
Resisc-45 18.9k 6.3k Remote Sensing
NAIP 244.4k 55.5k 200k Remote Sensing
EuroSAT 16.2k 5.4k Remote Sensing (Multi-spectral)
Camelyon17 302.4k 33.6k 200k Medical Imaging
iWildcam 129.8k 7.3k 150k Wildlife
GlobalWheat 2.9k 0.4k 80k Agricultural
VisDA2017 152.3k 55.4k 200k Synthetic-to-Real
Table 18: Dataset splits and sizes used in our experiments, as well as suggested number of ExPLoRA pre-training iterations.
@@ Appendix D Environmental Impact
Following (Cong et al., 2022), we compare the carbon footprint of pre-training using ExPLoRA with domain-specific solutions such as SatMAE. We use the carbon footprint calculator proposed by Lacoste et al. (2019). Our results are in table 19888Note: while the SatMAE paper reported 768 GPU hours for pre-training on fMoW-RGB and 576 GPU hours on fMoW-Sentinel, our measurements in Tables 1-5 of the main text have a revised number to ensure consistency of hardware platforms across comparisons.